OpenStudy (anonymous):

Find an equation of a line parallel to y=5x+5 and passing through the point (3, 2).

OpenStudy (anonymous):

The equation is in the form y=mx + b where the m is the slope of the line. A parallel line has the same slope as the equation above, so your new line needs to also have the slope 5. Using y=5x + b, plug in the coordinates (3,2) 2 = 5 (3) + b and solve for b. 2 = 15 + b b = -13 Therefore, your equation is y=5x - 13